Use a suitable identity to get each of the following products.
 (- a + c) (- a + c)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(- a + c) (- a + c)=(– a + c)^2
= c^2+ a^2–2ac
Using (a-b)^2= a^2+ b^2 - 2ab

Thanks